OpenStudy (haltiamali):

(-5x^7)(6x^7) QUICK ANSWER PLEASE!!

OpenStudy (jclark):

Follow this: x*x = x^2 x*x*x = x^3 x*x*x*x = (x^2)(x^2) 5x^2 = 5*x*x remember what symbols and placements imply multiplication. Remember what the exponent tells you to do with the base. It's all tied into multiplicaiton

OpenStudy (haltiamali):

-30x^49?

OpenStudy (jclark):

Closer. You did -5*6 correctly. What is the exponent 7 telling you to do to x?

OpenStudy (haltiamali):

Add them? -30x^14

OpenStudy (jclark):

there you have it

OpenStudy (haltiamali):

Can u help me with this one too? 8-b^3

OpenStudy (jclark):

(-5x^7)(6x^7) = (-5*x*x*x*x*x*x*x)(6*x*x*x*x*x*x*x) =-5*6*x*x*x*x*x*x*x*x*x*x*x*x*x*x =-30 times x multiplied by itself 14 times

OpenStudy (jclark):

same thing: 8 times -b^3. What does the exponent 3 telling you to do?

OpenStudy (haltiamali):

I need to do the 3rd root? right?

OpenStudy (haltiamali):

and thats 2

OpenStudy (jclark):

No, think about how you did the previous one. haltiamali: Add them? -30x^14

OpenStudy (jclark):

If it was asking for the third root, it would look like this: \[\sqrt[3]{8}\]

OpenStudy (haltiamali):

Ummm lol

OpenStudy (haltiamali):

idk what to do lol

OpenStudy (jclark):

wait, what is the problem asking you? what are the instructions? The second one you gave can't be reduced or simplified any further

OpenStudy (haltiamali):

Factor the polynomial

OpenStudy (jclark):

ooohhhh, my bad! This is 8 - b^3 \[8 -b ^{3}\]

OpenStudy (jclark):

What do you know so far? You mentioned third root... what do you know about the difference of cubes?